# AI-assisted Programming May Decrease the Productivity of Experienced Developers by Increasing Maintenance Burden

### v2

@@
We studied the code productivity increase introduced by Copilot by lines of code added, commits [^10] A commit is the fundamental unit of change on GitHub. Similar to saving a file that’s been edited, a commit records changes to one or more files on GitHub - https://docs.github.com/en/pull-requests/committing-changes-to-your-project/creating-and-editing-commits/about-commits and PRs submitted. Maintenance-related activities were assessed through PR reviews and PR rework, which measures the extent to which contributions initially submitted to the project need to be revised before they are integrated into the project. Based on our analysis of a large-scale panel dataset from GitHub, covering 2,755 projects and 1,699 contributors, we find that while AI adoption leads to productivity gains, they also increase maintenance-related activities due to a higher volume of review and rework needed per PR. Specifically, our analysis reveals a double-edged effect of GitHub Copilot on OSS development. At the project level, Copilot adoption is associated with a significant boost in productivity: projects that supported Copilot saw increases in lines of code added, commits, and PRs. However, this surge in contributions came also with an increase in PR rework (2.4% more code rework), indicating greater maintenance demands and possible declines in the quality of code initially submitted. At the contributor level, we observe an important redistribution of effort: less experienced, peripheral contributors increased their development activity, taking advantage of Copilot’s ability to lower coding barriers. Specifically, peripheral contributors, particularly those in the bottom percentiles in terms of experience, increased their commit activity by 43.5% and submitted 17.7% more PRs. In contrast, the core contributors reduced their commit activity by 19%, shifting their focus toward reviewing and maintaining code (a 6.5% increase), and shouldering a heavier quality assurance burden. Together, these findings highlight how AI can enable broader participation in OSS, but also raise concerns about the sustainability of these gains and the strain placed on a shrinking pool of experienced contributors who maintain quality in OSS projects.
1 Results
1.1 Project (Repository) Level Analysis
We conducted a DiD analysis using a panel dataset of OSS projects hosted on GitHub. The technical preview of GitHub Copilot was launched on June 29, 2021, providing a natural experiment for studying Copilot’s effect on code quality and maintainability. We focus on OSS projects owned by Microsoft, as the company had exclusive access to the technology behind GitHub Copilot, and therefore had greater access to the tool during its technical preview. [^11] In September 2020, Microsoft gained exclusive access to OpenAI’s GPT3. This access allowed Microsoft to repurpose and modify the model for code generation, leading to the development of GitHub Copilot. In June 2021, GitHub Copilot was launched as a technical preview. Anecdotal evidence and our interviews of Microsoft employees indicates that during the technical preview, access to Copilot was restricted to selected GitHub users, specifically employees of Microsoft/GitHub and maintainers of popular projects. This restricted access suggests that the primary users of Copilot during the technical preview were likely Microsoft and GitHub employees, along with selected Github uses who volunteered for the beta testing the software. [^12] Dog Fooding; is Microsoft speak for internal use of their own software, often to stress test its new software before public release - https://devblogs.microsoft.com/oldnewthing/20110802-00/?p=10003131313https://www.mckinsey.com/industries/technology-media-and-telecommunications/our-insights/thomas-dohmke-on-improving-engineering-experience-using-generative-ai We defined our observation period as the 12 months before and after this date, spanning from July 2020 to July 2022. During the technical preview of Copilot, it was endorsed for five programming languages — Python, JavaScript, Ruby, TypeScript, and Go. [^14] https://github.blog/news-insights/product-news/introducing-github-copilot-ai-pair-programmer/ At the project level, we define the treatment group as repositories whose primary programming language was among those endorsed by GitHub Copilot during the technical preview, while repositories using non-endorsed languages (e.g., C, C#, C++, Java, PHP, R, Scala) serve as the control group. Our project-level dataset consists of 2,755 repositories, with 1,660 in the Copilot-endorsed treatment group and 1,095 in the non-endorsed control group.
On GitHub, a PR is a vehicle of contribution through which contributors participate in the development process. A PR typically contains one or more commits and often represents a "patch" or feature addition to the project. Any individual can submit a PR (a request to merge their contribution into the project), which is then (peer) reviewed by the core contributors, who have write access to the source code. The core contributors reviewing the PR can decide to merge the PR into the main project, request modifications or, reject it. If modifications are requested for a PR, the author of the PR can address the comments and modification requests and re-submit the code in the form of follow-up commits for another round of review. This rework process continues until all issues with the code are resolved and the code can be merged, or until the idea behind the code is no longer in alignment with the project goals and the PR is rejected. Considering this, we can measure the amount of rework done on a PR submitted by a contributor by identifying the number of commits that are added to the PR after its initial submission. We use the PR rework effort measure as a proxy to determine the initial quality of code contributions submitted to the project and the maintenance-related efforts associated with the code contributions.
We analyzed the number of lines of code added, commits, and PRs as code development activities, and PR rework as maintenance related activities. Table 1 presents the estimations. After the introduction of GitHub Copilot, repositories whose primary programming language was endorsed by Copilot (treatment) experienced a significant increase in code development activities, such as the number of lines of code added ($\beta=0.163$, $SE=0.06$; $t=2.73$; $p<0.01$; $95\%$ CICI, $(0.046,0.28)$), commits ($\beta=0.04$, $SE=0.022$; $t=1.84$; $p<0.1$; $95\%$ CICI, (0.003,0.082)(-0.003,0.082)), and PRs ($\beta=0.042$, $SE=0.015$; $t=2.81$; $p<0.01$; $95\%$ CICI, $(0.013,0.071)$). This increase of productivity aligns with the analysis of industry experts (Peng et al. 2023) and the findings of recent academic research (Yeverechyahu et al. 2024).
Activity: Development Maintenance
DV: Lines of Added Code Commits Pull Requests PR Rework
Copilot 0.163*** 0.04* 0.042*** 0.024**
(0.06) (0.022) (0.015) (0.01)
Project FE
Month FE
PR Controls
N 66,120 66,120 66,168 66,168
Adj. $R^{2}$ 0.516 0.603 0.691 0.81
More importantly, we find that the submitted PRs requires greater amount of rework, indicating a likely decrease in the quality of the submitted PRs. When analyzing the PRs submitted, the increased amount of code resubmissions remains significant ($\beta=0.024$, $SE=0.01$; $t=2.35$; $p<0.05$; $95\%$ CICI, $(0.004,0.043)$), even when controlling for the number of PRs, indicating an increased demand for code reviews per unit of code development. Specifically, we find that treatment repositories experienced 2.4% more code rework, keeping the number of PR submitted constant. This increase likely reflects a higher volume of lower-quality code submissions to the project.
Figure 2 presents the event time analysis of the treatment effect on PR rework (leads-lags estimates). The coefficients for the pre-treatment periods are statistically insignificant, indicating that the parallel trends assumption holds (Angrist and Pischke 2009). The leads-lags estimate also indicate a long-term effect of Copilot. With greater deployment and adoption of AI, we expect increased contributor engagement and more frequent usage of Copilot. There is an upward trend in the post-treatment coefficients for PR rework that gradually become more positive over time (regression results are provided in the Supplementary Materials, Section 5.2). The observed pattern suggests a growing trend of increased code rework in repositories that supported AI coding partners.
Figure 2: Parallel trends and dynamic effects of the Copilot treatment for pull request rework. The horizontal axis represents the months relative to introduction of Copilot, while the vertical axis shows the estimated coefficients with confidence intervals (95%). The coefficients for the pre-treatment periods (leads) are statistically insignificant, indicating that there are no systematic differences in trends between the treated and control groups before the introducing of Copilot. This suggests that the parallel trends assumption holds, supporting the validity of our DiD estimation.
1.2 Contributor Level Analysis:
To understand how AI adoption affected contributors to OSS projects, we analyzed their contribution behavior before and after the introduction of GitHub Copilot. At the contributor level, we define treated contributors as those whose primary development activity is in Copilot-endorsed languages, while contributors working primarily in non-endorsed languages serve as the control group. Similar to our project level analysis, we use the technical preview of Copilot as our treatment period and perform DiD analysis at the contributor level. In total, we collected an individual level dataset consisting of 1,699 contributors from GitHub, with 1,186 in the Copilot-supported treatment group and 513 in the non-Copilot-supported control group. [^15] There are 37,334 contributors who made at least one contribution to the repositories in our sample. The contributions include, for example, posting a comment, submitting a commit, or conducting a pull request review. Among them, we filtered out 5,308 contributors who participated in more than three of the repositories we studied (we selected three repos to ensure sufficient variation for our PR reviewed repositories measure) . Then, we applied a programming language filter to construct a comparable treatment and control group for the individual-level analysis, and the data set was reduced to 1,699 contributors (who were users of the treated and control programming languages). Following the conceptualization of core contributors based on activity from past literature (Setia et al. 2012, Crowston et al. 2006), we used contributors’ level of activity during the pretreatment period, measured by the number of commits, to classify contributors into core contributors (top 25%) and peripheral contributors (rest 75%). From Figure 3, we observe that core contributors perform the majority of development activities (both in terms of commits and PR) in the projects.
We next examined commits and PRs as code development activities, and PR reviews (PR controlled) and PR reviewed repositories (PR controlled) as maintenance related activities. Based on our DiD analysis (Table 2), we find that core contributors engage in fewer development activities and in more maintenance activities after the launch of Copilot. The core contributors performed significant less code commits ($\beta=-0.357$, $SE=0.052$; t=6.83t=-6.83; $p<0.01$; $95\%$ CICI, (0.46,0.255)(-0.46,-0.255)) submissions. At the same time, the core contributors reviewed more PRs and do so across a broader set of repositories ($\beta=0.045$, $SE=0.018$; $t=2.42$; $p<0.05$; $95\%$ CICI, $(0.008,0.081)$). This shift suggests that core contributors are not only spending more of their limited time on maintenance tasks but also spreading themselves thinner across more repositories. The resulting burden of maintenance may come at the cost of reduced productivity among more experienced contributors.
To further examine the shift in workload across different contributor groups, we conducted a serious of subgroup DiD analyses based on the volume of contributions made by GitHub contributors. We calculated the quantity of commits submitted during the pretreatment period, and split the contributors accordingly into four percentile groups: 0 to 25%, 25% to 50%, 50% to 75% and 75% to 100%. The results are presented in Figure 4. The statistical results are presented in supplementary analysis (Section 4.2).
Table 2: The impact of technical preview on GitHub’s core contributor behaviour.
Activity: Development Maintenance
DV: Commits Pull Request PR Review PR Reviewed Repos
Copilot 0.142*** 0.091*** 0.04 0.008
(0.408) (0.03) (0.023) (0.012)
Core Contributor -0.357*** -0.027 0.06* 0.045**
$\times$ Copilot (0.052) (0.042) (0.035) (0.018)
Individual FE
Month FE
PR Controlled
N 51,408 51,408 51,408 51,408
Adj. $R^{2}$ 0.643 0.607 0.799 0.757
(a) Commits.
(b) Pull Request.
(c) PR Reviews.
(d) PR Reviewed Repos.
Figure 4: Contribution activities analysis by contributor subgroup. Panels show estimated coefficients (converted to %) from DiD regressions with 95% confidence intervals, capturing the relative change in activity post-Copilot exposure compared to control repositories. (a) Commits: Conversely, commit activity declines progressively with contributor experience, with the top 25% experiencing a 19% reduction, suggesting reduced hands-on coding engagement. (b) Pull Requests: The most peripheral contributors (0–25%) significantly increase their PR submissions (17.7%), highlighting increased participation from less experienced developers. (c) PR Reviews: The top 25% of contributors (core) exhibit a significant increase in review activity, suggesting a shift of responsibility towards the core contributors. (d) PR Reviewed Repositories: Similarly, only the core contributor group shows a meaningful rise in the number of distinct repositories reviewed, indicating a broader oversight role.
2 Discussion
Recent research suggests that the adoption of GenAI technologies brings not only productivity gains but also a host of unintended secondary effects. These include anchoring bias (Chen and Chan 2024), reinforcement of societal and algorithmic biases (Williams-Ceci et al. 2025, Nicoletti and Bass 2023), reduced collective diversity (Doshi and Hauser 2024), dialect prejudice (Hofmann et al. 2024), and snowball effects that amplify initial biases over time (Glickman and Sharot 2025). Meta-analyses further show that in decision-making tasks, human-AI teams can perform worse than either alone, despite outperforming in content generation tasks (Vaccaro et al. 2024).
In software development communities like OSS, the key concern is how AI affects code quality and software security. By lowering the skill barrier for writing code (Dakhel et al. 2023), AI tools can encourage broader participation but may also lead developers to rely on AI-generated code without fully understanding its implications (Barrett et al. 2023). This growing dependency increases the risk of bugs, security vulnerabilities, and fragile or suboptimal code being introduced into production. In the face of unprecedented demand for OSS infrastructure (Nagle 2019), the cost of such quality-related issues can be severe. As a case in point, in 2011 a major vulnerability (nicknamed Heartbleed) in an OpenSSL OSS project was overlooked and included in an update of the project. It went unnoticed for years, allowing any sophisticated hacker to capture secure information being passed to vulnerable web servers, including passwords, credit card information, and other sensitive data. The widely held view on Heartbleed underscores the risks of under-resourced maintenance related activities in OSS projects: : [^16] https://mashable.com/archive/heartbleed-bug-websites-affected
Our study sheds light on the hidden maintenance burdens AI introduces into OSS development communities. While AI-assisted coding lowers entry barriers and attracts peripheral (often less experienced) contributors, potentially fueling innovation, it also introduces substantial maintenance challenges. Chief among these are longer rework cycles and declining code quality, which heighten the risk of future security issues. More critically, the burden of reviewing and correcting AI-generated code increasingly falls on experienced core contributors diverting their attention from creative tasks like writing new code to routine and reactive maintenance work such as reviewing and reworking others’ contributions (Eghbal 2020, 2016).
In sum, this research contributes to our understanding of AI driven programming tools in three key ways. First, we find that the introduction of Copilot increased programmer productivity, resulting in a higher volume of code contributions, measured through commits and PRs, at the repository level. At the individual level, OSS contributors submit more PRs after the adoption of GitHub Copilot. This finding aligns with the industry study conducted by Microsoft (Peng et al. 2023). Second, and more importantly, we provide empirical evidence illustrating how AI adoption can also subsequently increase the need for maintenance related activities in OSS communities. We find that the code submissions (PRs) after the introduction of AI require more rework. This increase in rework adds complexity to the review process before PRs can be merged, potentially undermining the project’s quality. Poor-quality code can result in more bugs, increased rework, and eventual large-scale refactoring of the entire project (Barrett et al. 2023). Lastly, we examine the heterogeneous effects of AI adoption across different groups of contributors. Our results show that core contributors, not only review more PRs while contributing fewer commits, but also extend their maintenance related activities across a wider range of repositories. This suggests that AI-assisted code submissions from less experienced developers may shift the burden onto core maintainers, ultimately undermining their productivity. As these key contributors become overextended, it may increase the risk of unresolved bugs and security vulnerabilities, threatening the long-term sustainability of OSS projects.
To illustrate the scale of Copilot’s impact on OSS communities, Microsoft core contributors in our dataset conduct on average, 976 commits, 160 PRs, and 166 PR reviews annually before its introduction. The increased volume of code associated with Copilot adoption results in an additional workload – each core contributor is expected to review approximately 10 more pull requests annually. This added maintenance burden corresponds to a reduction of 164 commits and 9 pull request contributions per year per core contributor. More critically, GitHub’s 2024 surveys reveal that more than one-third of contributors to the 10 most popular OSS projects made their first contribution after signing up for GitHub Copilot, highlighting a significant influx of new and often less experienced developers [^17] https://github.blog/news-insights/octoverse/octoverse-2024/. With annual contributions to OSS projects approaching 1 billion, this surge in participation significantly increases the burden on core contributors, who take on the maintenance related tasks in the project. As a result, maintainers are compelled to reallocate their time toward reviewing and managing code submissions instead of writing new code.
Prior beliefs have largely emphasized the productivity and efficiency gains of AI pair programming, fueling expectations that tools like Copilot would significantly accelerate software development (Peng et al. 2023). While these benefits are real, our results reveal a more complex picture. Specifically, we find that Copilot also amplifies software maintenance challenges—particularly for core contributors—as evidenced by our individual-level analysis. The reliance on AI tools may also impact the foundational learning of peripheral developers. With AI providing quick solutions, there is a risk that developers may not fully engage with the underlying principles of coding nor the best coding practices, leading to a superficial understanding. This gap in comprehension can result in code that is functional but lacks robustness, making future maintenance more challenging. This concern aligns with findings from AI-assisted customer support settings, where less experienced workers see large productivity gains, while more skilled workers experience only modest improvements (Brynjolfsson et al. 2025).
The implications of AI adoption in OSS communities extend well beyond the boundaries of OSS development. As OSS practices and tools are increasingly integrated into corporate and institutional workflows, the quality, reliability, and sustainability challenges observed in OSS are likely to appear in other areas of software development. Many firms now rely heavily on OSS libraries, frameworks, and platforms, and increasingly adopt the open-source approach to develop software and products (Nagle 2019). As a result, any decline in code quality or strain on maintenance capacity within OSS can cascade into broader software ecosystems, affecting enterprise systems, digital services, and public-sector infrastructure that depend on these components.
As a concluding remark, the challenges observed in OSS, such as quality concerns and increased maintenance burdens driven by productivity gains among lower-skilled contributors, should serve as an early warning for similar risks in other domains where AI is being promoted to boost productivity and innovation.
3 Methods

### v3

@@ 1 Introduction
Based on our analysis of a large-scale panel dataset from GitHub, we find that while AI adoption leads to productivity gains, they also increase maintenance-related activities due to a higher volume of review and rework needed per PR. Specifically, our analysis reveals a double-edged effect of GitHub Copilot on OSS development. At the project level, Copilot adoption is associated with a significant boost in productivity: projects that supported Copilot saw increases in lines of code added, commits, and PRs. However, this surge in contributions came also with an increase in PR rework (2.4% more code revisions), indicating a possible decline in the quality of code initially submitted resulting in an accumulation of technical debt. At the contributor level, we observe an important redistribution of effort: the peripheral contributors (the less active contributors to the projects) increased their development activity, taking advantage of Copilot’s ability to lower coding barriers. Specifically, peripheral contributors, particularly those in the bottom percentiles in terms of their previous contributions, increased their commit activity by 43.5% and submitted 17.7% more PRs. In contrast, the core contributors reduced their commit activity by 19%, shifting their focus toward reviewing and maintaining code (a 6.5% increase), and shouldering a heavier quality assurance burden. Together, these findings highlight how AI can enable broader participation in OSS, but also raise concerns about the sustainability of these gains and the strain placed on a shrinking pool of experienced contributors who maintain quality in OSS projects.
2 Related Literature and Conceptual Development
2.1 AI Assisted Code Development in OSS Projects
The rapid advancement of GenAI technologies, particularly LLMs like ChatGPT and GitHub Copilot, is transforming how software is developed and how online knowledge communities operate. Current research on AI-assisted code development has shown the substantial impact of the technology on productivity. A study to assess the productivity benefits of using Copilot revealed that developers who use Copilot completed their programming task 55.8% faster than the control group (Peng et al. 2023). Another study by GitHub reports that the use of Copilot Chat increases programmers’ confidence, with participants self-reporting improvements in code readability, reusability, conciseness, maintainability, and resilience (Rodriguez 2023). These productivity gains also translate into labor-market outcomes: developers exposed to AI-assisted coding experience faster career progression in the short- to medium-term (Li et al. 2025). Work also finds that AI coding assistants reshape the allocation of work. For instance, Yeverechyahu et al. (2024) investigate the impact of GitHub Copilot on innovation in OSS projects. They find a significant increase in overall code contributions, accompanied by a shift in the nature of innovation toward more routine and incremental changes. Song et al. (2024) find that Copilot adoption increases project-level code contributions, though this comes at the cost of an increase in coordination time for code integration. Relatedly, Hoffmann et al. (2025) show that access to GitHub Copilot reallocates developers’ effort toward core coding tasks and away from project management and coordination activities.
While AI-assisted code development promises substantial productivity gains, its implications for software maintenance remain less well understood. Prior research on software development has long recognized that development costs are often small relative to maintenance costs, which include sustaining activities associated with ensuring software quality and security (Nagle 2019). In the case of OSS, while users can benefit from reduced up-front costs, collective intelligence of the crowd, and flexibility to implement changes, the challenges of maintenance get magnified as contributors are not contractually obligated to maintain the software (von Hippel and von Krogh 2003, Nagle 2019). The Linux Foundation’s OSS Contributor Survey provides insightful perspectives on the complexities involved in maintaining OSS (Nagle et al. 2020). Firstly, it highlights that “general housekeeping" tasks, such as project maintenance, bug reporting and documentation, and organizational or administrative duties, often consume a more significant portion of contributors’ time than desired. Secondly, despite a preference among contributors to spend less time on maintenance tasks, there’s a broad acknowledgment of the importance of these activities, especially those related to software security, for the success and integrity of their projects (Nagle et al. 2020).
Furthermore, AI code assistants, including prompt-based and “vibe coding” practices, promise to increase productivity while easing access for contributors to submit code, even in complex and mature OSS projects. Recent work has begun to examine vibe coding as an emerging and controversial paradigm in AI-assisted software development, in which programmers rely on natural language interaction with generative models to maintain flow and rapidly explore solutions, often with minimal upfront specification (Pimenova et al. 2025, Fawzy et al. 2025). While this approach can substantially accelerate development and foster experimentation, the literature consistently highlights associated risks, including underspecified requirements, reduced reliability, difficulties in debugging, increased latency, and heavier burdens on code review and collaboration (He et al. 2025). A recurring theme is a speed–quality paradox (Fawzy et al. 2025): although vibe coding enables rapid progress and is often perceived as highly efficient by its users, the resulting code is frequently described as fast but flawed. Moreover, quality assurance practices are commonly deprioritized in vibe coding workflows, with users skipping testing, accepting AI-generated outputs with little modification, or delegating verification back to the AI itself. These patterns raise concerns about latent defects and heightened maintenance demands, particularly in collaborative and production-oriented software environments (Schreiber and Tippe 2025).
These maintenance-related concerns are amplified by evidence that less experienced programmers appear to benefit more from the use of Copilot. For novice developers, AI-driven code suggestions can help them overcome steep learning curves and contribute even to mature OSS projects. However, both Dakhel et al. (2023) and Peng et al. (2023) caution that novice programmers may place undue trust in AI-generated code without fully understanding its broader design or integration implications, increasing the likelihood of downstream quality and integration challenges. Over time, these issues can accumulate as technical debt, increasing the maintenance burden for OSS projects and shifting responsibility toward a small group of experienced core contributors. Anecdotal evidence [^7] https://stackoverflow.blog/2025/08/07/a-new-worst-coder-has-entered-the-chat-vibe-coding-without-code-knowledge/ suggests that while AI-assisted coding can produce seemingly complete solutions with minimal effort, these outputs frequently exhibit substantive deficiencies when reviewed by experienced developers. Furthermore, Schreiber and Tippe (2025) show that even when AI-generated code passes functional tests, it often embeds latent security weaknesses and maintainability issues that shift effort from development to downstream debugging and remediation, effectively accumulating technical debt. Relatedly, the growing volume of low-quality, AI-generated security artifacts has begun to strain OSS security workflows. In January 2026, the cURL open-source project discontinued its bug bounty program after being overwhelmed by a flood of AI-generated vulnerability reports that consumed reviewer attention without yielding actionable findings  [^8] https://www.techradar.com/pro/security/curl-will-stop-bug-bounties-program-due-to-avalanche-of-ai-slop. In other words, the ease of generating code masks underlying structural and logical gaps that become apparent under expert scrutiny (Barrett et al. 2023). As with other AI tools that create superficially polished but potentially shallow results, vibe coding may be appealing for rapid prototyping, yet the secondary effects that this may have on reliability or maintainability when applied to nontrivial software development tasks is an open question.
Our study investigates this issue by analyzing how the adoption of GitHub Copilot reshapes contributor effort across production and maintenance activities, and whether it disproportionately shifts the burden of maintenance onto a smaller group of core contributors.
2.2 Conceptualization of Technical Debt
Technical debt captures the intertemporal trade-off between short-term development speed and long-term system maintainability (Cunningham 1992). A growing body of research has established that technical debt is economically consequential, persistent, and shaped by organizational choices. Drawing on the information systems and computer science literature, we synthesize prior work in Table 1 and adopt a definition of technical debt in this study as accumulation of suboptimal code, design, or documentation decisions in PR process that increase PR Rework (technical debt) and PR Review (maintenance effort) - a definition that is particularly prominent in the era of GenAI software development.
Measurement approaches for technical debt have evolved along two complementary axes. One strand emphasizes static code metrics (e.g., code complexity, duplication, test coverage) as proxies for debt accumulation, enabling automated detection and cross-project comparisons (Paramitha and Massacci 2023, Yoo et al. 2025). Paramitha and Massacci (2023) show how dependency structures in open-source ecosystems (top 600 Python packages) amplify technical leverage, allowing localized weaknesses to propagate across packages. Yoo et al. (2025) similarly demonstrate that dependency network structures shape the diffusion of security vulnerabilities in software supply chains. Another strand operationalizes technical debt through process and outcome indicators—such as rework rates, defect density, time-to-merge, and maintenance effort—thereby connecting debt to developer behavior and organizational performance (Ramasubbu and Kemerer 2016, Banker et al. 2021, Ramasubbu and Kemerer 2021). Both approaches have value: the former metric-based techniques are scalable and actionable for tooling, while the later process indicators capture the economic and human costs that managers and maintainers ultimately face.
Empirical studies across industry and OSS ecosystems consistently find that technical debt impairs long-run productivity and increases maintenance burdens (Paramitha and Massacci 2023). Work in empirical software engineering shows that higher measured debt correlates with increased bug rates, longer defect resolution times, and reduced velocity for feature delivery (Ramasubbu and Kemerer 2016). Rinta-Kahila et al. (2023) further demonstrate that organizations can become “trapped” in technical debt due to coordination failures, organizational inertia, and escalating switching costs. The introduction of GenAI tools such as GitHub Copilot represents a departure from these assumptions (Yeverechyahu et al. 2024). Unlike prior productivity-enhancing tools, GenAI dramatically lowers the marginal cost of producing code and reduces skill barriers to contribution (Peng et al. 2023). While prior studies document productivity gains from AI-assisted coding, they provide limited insight into how these gains translate into technical debt and maintenance related challenges.
From a technical debt perspective, GenAI may accelerate debt accumulation by increasing code volume without proportionate improvements in quality, the integration with the software, or architectural coherence (Pimenova et al. 2025). By lowering the cost of producing code, AI-assisted programming encourages rapid iteration and experimentation, but can shift attention away from longer-term concerns such as maintainability, readability, and alignment with existing system design (Barrett et al. 2023, Schreiber and Tippe 2025). As a result, defects, inconsistencies, and design shortcuts may be introduced more quickly than they can be identified and resolved. Over time, this imbalance can compound, transforming short-term productivity gains into persistent maintenance obligations that must be absorbed by experienced developers (Eghbal 2020). While GenAI promises to enhance development speed and broaden participation, it may simultaneously intensify the very technical debt that constrains system reliability, scalability, and long-run performance (Ramasubbu and Kemerer 2021). If this logic is valid, we expect that GenAI may lead to a greater accumulation of technical debt in OSS projects.
Table 1: Selected Literature on Technical Debt and Software Maintenance
Study Context Method Measurement Key Findings
Banker et al. (2021) Customer relationship management (CRM) systems in 26 firms Econometric analysis Percentage of customized codes in the CRM system that do not adhere to vendor-prescribed standard Higher technical debt is associated with lower firm performance and reduced operational efficiency over time.
Ramasubbu & Kemerer (2021) Outsourced Commercial Off-The-Shelf (COT) enterprise systems Econometric analysis Violations of the design and programming standards established by the vendor of the COTS enterprise system Active remediation policies reduce long-term maintenance costs, but excessive deferral leads to escalating technical debt.
Ramasubbu & Kemerer (2016) Enterprise software systems Econometric analysis Client customizations violating vendor standards (business logic / data schema), API violations Accumulated technical debt increases system failure hazards and shortens system lifespan.
Paramitha & Massacci (2023) Top 600 Python packages, Python open-source ecosystem Box plot, simulation Technical leverage as the ratio between dependencies (other people’s code) and own codes of a software package Technical leverage amplifies maintenance risks through dependency networks, increasing downstream technical debt.
Yoo et al. (2025) Software supply chains Econometric analysis Dependency centrality, vulnerability exposure Complex dependency structures increase vulnerability propagation, linking technical debt to security risks.
Rinta-Kahila et al. (2023) Legacy system replacement Qualitative case study, system-dynamics model Sociotechnical debt indicators: architectural compromises, technical inertia & digital options foregone or constrained Technical debt becomes entrenched through organizational routines, creating lock-in and delaying modernization.
Rolland et al. (2018) Digital platforms in firms Longitudinal case analysis The interactions between digital options and digital debt as a contribution to understanding the complex choices organization face in managing digital platforms Strategic flexibility often trades off with accumulating digital and technical debt.
This study OSS Repositories on GitHub Econometric analysis Accumulation of suboptimal code, design, or documentation decisions in PR process that increase PR Rework (technical debt) and PR Review (maintenance effort) Code written after the adoption of AI requires more rework to satisfy repository standards
2.3 Technical Debt in OSS projects
Prior research on software quality and technical debt has largely focused on proprietary software developed within organizational boundaries, where development and maintenance are often treated as sequential and centrally coordinated activities (Paramitha and Massacci 2023, Yoo et al. 2025, Ramasubbu and Kemerer 2016, Banker et al. 2021, Ramasubbu and Kemerer 2021). In contrast, OSS is produced through a distributed development model in which participation is open and contributions can be submitted by anyone, blurring the boundary between development and maintenance and creating distinct challenges for managing quality and technical debt (Nagle 2019). Therefore, our study focuses on the PR review process, a core coordination mechanism in continuous integration and continuous deployment (CI/CD) of software.
The development workflow of an OSS project is illustrated in Figure 1. A feature of this development workflow is that it allows multiple contributors to work independently on separate branches (or forks) and submit their changes for inclusion in the main project branch. This is done through a PR, which serves as a formal request to review and merge code contributions (such as new features or improvements) made by project contributors. Core or experienced contributors often handle key maintenance tasks: they review submitted PRs for quality and compliance, suggest improvements or corrections adhere to project standards (Banker et al. 2021), and integrate approved changes into the main project. This workflow enables distributed development and community-driven innovation, allowing peripheral contributors who come from the community of users of the project to contribute. At the same time, it ensures that the code is reviewed, refined, and improved before merging into the main branch. The effectiveness of this review process has helped OSS achieve remarkably high quality, surpassing proprietary software in metrics like bugs per 1,000 lines of code. [^9] https://blogs.worldbank.org/en/opendata/quality-open-source-software-how-many-eyes-are-enough However, the success and effectiveness of this "decentralized-development" approach hinges on the core contributors who actively participate in the review-rework process of the PRs submitted by different contributors (Rullani and Haefliger 2013).
Figure 1: The workflow of OSS projects. It comprises of a primary branch of a GitHub (project), which typically contains the main source code that serves as the foundation for new feature development, bug fixes, and updates. Changes to this branch are usually controlled through a structured review process conducted by maintainers to ensure code quality and prevent issues.To contribute to a project, contributors submit a PR to merge their code changes, such as new features or improvements. It allows developers to submit modifications, request feedback, and merge updates into the main branch. The review and revision process comprises of two activities that we measure in our study - PR review and PR rework. Eventually the reviewed and reworked changes will either be merged into the main branch or undergo additional rounds of review and revision.
Focusing on the PR review process is particularly well suited to OSS contexts, where code is continuously developed, reviewed, and deployed into production, often with rapid and frequent version updates (Roberts et al. 2006, Medappa et al. 2023). By analyzing PR reviews and rework as they occur, our study captures maintenance activities in real time, providing a more accurate representation of how quality control and technical debt are managed within distributed development workflows. Importantly, the PR review process allows us to observe technical debt at the point of entry—when new code is evaluated, revised, or rejected prior to integration. This technical debt measurement would be applied to other CI/CD software projects beyond OSS community.
Prior work conceptualizes technical debt as deferred maintenance or suboptimal design choices that increase future rework costs, typically operationalized through downstream quality degradation or maintenance outcomes in traditional, organization-bound software development settings (Ramasubbu and Kemerer 2016, Banker et al. 2021, Ramasubbu and Kemerer 2021). In contrast, OSS development unfolds through continuous integration and deployment of new features, where design trade-offs and quality concerns are often raised and resolved at the PR level. In this context, PR rework provides a direct operationalization of technical debt, capturing realized remediation effort during code integration.
3 Data and Methods